A tank with 2 m3 capacity initially contains 1 m3 liquid at 50 °C. There are two inlet streams to the tank and one outlet stream. The first inlet stream is steady at 0.2 m3 /min with T= 50 °C. Second inlet stream has a temperature of 100 °C and surges on a 2 min cycle. During the first minute the flow rate is steady at 0.2 m3 /min and during the second minute its flow stops completely, then the cycle repeats itself. The flow rate of the outlet stream is constant at 0.3 m3 /min. What will be the temperature in the tank at the end of the 4 minutes from the start of the operation (List the assumptions used)?
Answer #1

Doudm? Inin looc: © Sd Consider System as some Shown in which stream o enters at flowrete ord m/min (Lonste) n Strom ② entereT-63.74 + cé 0.36€ At to , T=50c. 50-60 pceo. C=3-44 : T= 6 7 g +87-14 e 036t At t=lmin , AS292127588 6 66 T = 63-11-13:lle.Vto, dv=-0.1 m Now, between t = 1- I min . Valmo From a -0-317-50) -01 -011 dt d1+0=27=15 dt Test feest osat+C. Teoost 75 e twom between ta3-4mm Veo, dv 1 2 0 m from min at (1-2 min ) v=lm? Solution will be some es betree T= 75+ceast At t= 3mn T-59:3

using given conditions at inlet at assuming no time delay and constant fluid properties we can find the temeprature at end of 4 min

for any calrification on solution please write in comments